# Quantum Spin Liquids in Weak Mott Insulators with a Spin-Orbit Coupling

### v1

@@ IV.2 Extent of the chiral spin liquid phase
[…]
@@
Figure 5: (a) One of the four rhombuses which has an edge $\langle a,b\rangle$ parallel to the bond $\langle i,j\rangle_{\mathbf{\delta}_{1}}$, oriented along $\mathbf{\delta}_{1}$, is shown in red. The rhombus in green has second nearest neighbour sites $(c,d)$ and has the bond $\langle i,j\rangle_{\mathbf{\delta}_{1}}$ as its diagonal. (b) One of the four rhombuses with the bond $\langle a,b\rangle_{-\mathbf{\delta}_{1}}$ oriented along $-\mathbf{\delta}_{1}$ that also has the bond $\langle i,j\rangle$ as a parallel edge. (c) A rhombus with the bond $\langle c,d\rangle_{\mathbf{\delta}_{1}}$ as its diagonal which has the second-nearest neighbour sites $(c,d)$.
@@ IV.3 Inducing a chiral spin liquid starting from a valence bond solid phase
$\left(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right){S}_{k}^{[x}{S}_{l}^{y]}\rightarrow\left(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right)\langle{S}_{k}^{[x}{S}_{l}^{y]}\rangle,$ (18)
[…]
@@ Appendix A Details of the Canonical Transformation
$T^{+}=$ $\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}n_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}h_{j,-\beta},$ (25)
$T^{-}=$ $\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}h_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}n_{j,-\beta},$ (26)
$T^{0}=$ $\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}\left(h_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}h_{j,-\beta}+n_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}n_{j,-\beta}\right),$ (27)
@@
$S$ is calculated order by order in $\tilde{t}/U$ so that fluctuations to the doubly occupied sector are eliminated at each order in $\tilde{t}/U$ We can expand the generator $S$, as $S=S^{(1)}+S^{(2)}+S^{(3)}+\cdots$, where $S^{n}\propto\left(\tilde{t}/U\right)^{n}$. To compute ring-exchange type terms we need to calculate $H_{\text{eff}}$ to $\mathcal{O}(\tilde{t}^{4}/U^{3})$. Therefore we require the expansion of the generator $S$ to $\mathcal{O}(\tilde{t}^{3}/U^{3})$ or up to $S^{(3)}$. We find that the first three terms of the expansion are:
@@
$iS^{(2)}=$ $\frac{1}{U^{2}}\left([T^{+},T^{0}]+[T^{-},T^{0}]\right)$ (32)
$iS^{(3)}=$ $\frac{1}{U^{3}}([[T^{+},T^{0}],T^{0}]-[[T^{-},T^{0}],T^{0}]),$
@@ Appendix B Full model with Spin-Orbit interactions
[…]
@@
[…]
@@ Appendix D DMRG Data
Using iDMRG we obtain a set of useful data that is used to characterize the nature of the quantum ground state. Firstly, iDMRG gives the ground state energy per lattice site. In addition to local magnetization $\langle S^{z}_{i}\rangle$, we obtain long-range real space spin correlations, $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle$, along the infinite direction $x$ of the cylinder. From the real space spin correlations, we can calculate the static spin structure factor (10). We also obtain real space long-ranged dimer correlations of the spins, $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle$, where $D^{n}_{i}=\mathbf{S}_{i}\cdot\mathbf{S}_{i+\mathbf{a}_{n}}$ is the dimer operator along the direction $\mathbf{a}_{n}$ (see Fig. 1 (a)). Using this the static dimer structure factor (12) can be calculated. Further, we also obtain the scalar chirality $\chi$ as defined in (9), and the handedness $\omega$ as defined in (11). Most importantly, from iDMRG we can obtain the momentum-resolved entanglement spectrum. Figs. 8-12 show the various correlation functions and entanglement spectra for all the different phases appearing in Figs. 2 & 3.
@@
Figure 8: Observables in the CSL phase ($J_{1}=1$, $J_{2}=0.01$, $J_{3}=0.01$, $J_{r}=0.2$, $D_{z}=J_{r_{1}}=0$). The real space spin correlations $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle_{c}=\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle$ are shown in (a), the static spin structure factor $S(\mathbf{k})$ in (b). Real space dimer correlations $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle_{c}=\left\langle D^{n}_{i}D^{n}_{j}\right\rangle-\left\langle D^{n}_{i}\right\rangle\left\langle D^{n}_{j}\right\rangle$ along the the different directions $\mathbf{a}_{n}$ are shown in (d)-(f), and the static dimer structure factors $D_{n}(\mathbf{k})$ in (g)-(i). Here we have chosen $\mathbf{R}_{0}=3\mathbf{a}_{1}$. The entanglement spectrum in (c) shows the characteristic degeneracy pattern of the Kalmeyer-Laughlin type CSL [66, 67].
@@
Figure 12: Observables in the magnetic ordered (MO) phase ($J_{1}=1$, $J_{2}=0.012$, $J_{3}=0.012$, $J_{r}=0.24$, $D_{z}=0.065$, $J_{r_{1}}=0.072$). The real space spin correlations $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle_{c}=\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle$ are shown in (a), the static spin structure factor $S(\mathbf{k})$ in (b). Real space dimer correlations $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle_{c}=\left\langle D^{n}_{i}D^{n}_{j}\right\rangle-\left\langle D^{n}_{i}\right\rangle\left\langle D^{n}_{j}\right\rangle$ along the the different directions $\mathbf{a}_{n}$ are shown in (d)-(f), and the static dimer structure factors $D_{n}(\mathbf{k})$ in (g)-(i). Here we have chosen $\mathbf{R}_{0}=3\mathbf{a}_{1}$. Even though the spin correlations in (a) are long-ranged, they are shorter-ranged than the $120^{\circ}_{\pm}$ phases in Figs. 10, 11.
@@ Appendix E Different types of $120^{\circ}$ order
To test this hypothesis, we have computed the overlap in the wavefunctions between representative states of these phases. For the representative state |$\ket{\psi_{120^{\circ}}}$ for the $120^{\circ}$ phase, we have used the ground state corresponding to $J_{1}=1$, $J_{2}=J_{3}=0.0$, $J_{r}=0.16$. For the |$\ket{\psi_{120^{\circ}_{+}}}$ state the corresponding parameters are $J_{1}=1$, $J_{2}=J_{3}=0.01$, $J_{r}=0.2$, $D_{z}=0.005$, $J_{r_{1}}=1.2$, and for the |$\ket{\psi_{120^{\circ}_{-}}}$ state they are $J_{1}=1$, $J_{2}=J_{3}=0.01$, $J_{r}=0.2$, $D_{z}=0.195$, $J_{r_{1}}=1.2$. We find that $|\langle\psi_{120^{\circ}}|\psi_{120^{\circ}_{+}}\rangle|\sim 0.7$ and $|\langle\psi_{120^{\circ}}|\psi_{120^{\circ}_{-}}\rangle|\sim 0.7$. Therefore to leading order the $120^{\circ}$ phase is an equal weight superposition of the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phases. Also we find, $|\langle\psi_{120^{\circ}_{+}}|\psi_{120^{\circ}_{-}}\rangle|\sim 0.05$, which suggests that there is minimal overlap between the two oppositely handed $120^{\circ}$ phases.
Further, to distinguish between the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phases, we use these representative states, |$\ket{\psi_{a}}$ ($a\in\{120^{\circ}_{+},120^{\circ}_{-}\}$), to compute the expectation value of the Hamiltonian parameterized along the axis $J_{r_{1}}/J_{r}=1.7$ of the phase diagram in Fig. 2 (a). We compute $\langle\psi_{a}|H|\psi_{a}\rangle$ as a function of $D_{z}/J_{1}$ and this is shown in Fig. 13. Even though the representative states are not eigenstates of the Hamiltonian, we observe in Fig. 13, a crossing in the values of $\langle H\rangle$ computed using the |$\ket{\psi_{120^{\circ}_{+}}}$ and |$\ket{\psi_{120^{\circ}_{-}}}$. This suggests that the two differently-handed $120^{\circ}$ phases are preferred on opposite sides of the crossing. The location of this crossing of $\langle H\rangle$ in Fig. 13, also roughly corresponds to the location of the phase boundary between the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ (see Fig. 2 (a)). The crossing in the value of $\langle H\rangle$ computed using representative states is similar to what we typically expect to see in a variational calculation using metastable eigenstates, however in Fig. 13 we do not perform a variational calculation.
@@
Figure 13: The expectation value of the Hamiltonian, $\braket{\psi_{a}}{H}{\psi_{a}}$ as a function of $D_{z}/J_{1}$, computed using representative states |$\ket{\psi_{a}}$ from the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phases (see Appendix E for details) along the $J_{r_{1}}/J_{r}=1.7$ axis in Fig. 2 (a). The crossing in the value of $\langle H\rangle$ between the $120^{\circ}_{+}$ and $120^{\circ}_{-}$ phases is at $D_{z}/J_{1}\approx 0.155$. This value is consistent with the location of the phase boundary in Fig. 2 (a), between these phases. The two types of oppositely handed $120^{\circ}$ order are visualized using classical spins in (b) and (c) for the $120^{\circ}_{+}$ ($\omega>0$) and $120^{\circ}_{-}$ ($\omega<0$) phases respectively. Along the red lines from bottom to top, (which is also a positive direction of $\mathbf{v}$) the spins in (b) and (c) wind anti-clockwise and clockwise respectively, indicating their handedness $\omega$.

### v2

@@ IV.2 Extent of the chiral spin liquid phase
In Fig. 2 (a), we find that the CSL phase is stabilized until $D_{z}/J_{1}=0.13,J_{r_{1}}/J_{r}=1.4$, even though $D_{z}/J_{r_{1}}\sim 0.5$ (the compensation condition discussed in Sec. IV.1). This indicates a failure of the compensation mechanism beyond a certain value of $J_{r_{1}}$. To understand the reason for the failure, we construct a heuristic Hamiltonian for the CSL phase, using the DM and the SOC-mediated ring exchange interactions. We decouple the $H_{J_{r_{1}}}$ term the same way as in Eq. (13), where now the expectation value $\langle\cdot\rangle$ is computed with respect to an iDMRG wavefunction from CSL phase. After the decoupling we can construct an effective DM type interaction, containing both the original DM term as well as the correction from $H_{J_{r_{1}}}$. Conceptually, we are mapping the two independent variables $D_{z}$ and $J_{r_{1}}$ of the phase diagram in Fig. 2 (a) into one variable $D_{z}$. This makes sense for the CSL phase since it appears in a narrow elongated region (with $D_{z}/J_{r_{1}}\sim 0.5$) which can be approximated as being one-dimensional for the purpose of this discussion. The effective Hamiltonian is given by:
@@
Figure 6: (a) One of the four rhombuses which has an edge $\langle a,b\rangle$ parallel to the bond $\langle i,j\rangle_{\mathbf{\delta}_{1}}$, oriented along $\mathbf{\delta}_{1}$, is shown in red. The rhombus in green has second nearest neighbour sites $(c,d)$ and has the bond $\langle i,j\rangle_{\mathbf{\delta}_{1}}$ as its diagonal. (b) One of the four rhombuses with the bond $\langle a,b\rangle_{-\mathbf{\delta}_{1}}$ oriented along $\mathbf{\delta}_{1}$ that also has the bond $\langle i,j\rangle$ as a parallel edge. (c) A rhombus with the bond $\langle c,d\rangle_{\mathbf{\delta}_{1}}$ as its diagonal which has the second-nearest neighbour sites $(c,d)$.
@@ IV.3 Inducing a chiral spin liquid starting from a valence bond solid phase
$\hskip-5.12128pt\left(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right){S}_{k}^{[x}{S}_{l}^{y]}\rightarrow\left(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j}\right)\langle{S}_{k}^{[x}{S}_{l}^{y]}\rangle,$ (18)
where the expectation value $\langle\cdot\rangle$ is computed in the ground state using the iDMRG wavefunctions at these particular parameters (this is different from Sec. IV.2, where an average over different points in the phase diagram was taken). Next, we compute $\Delta_{1,\mathbf{\delta}_{n}}$ and $\Delta_{2,\mathbf{\delta}_{n}}$ for each of the three types of bonds $\mathbf{\delta}_{n}$ in the triangular lattice:
@@ Appendix A Details of the Canonical Transformation
$T^{+}=$ $-\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}n_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}h_{j,-\beta},$ (25)
$T^{-}=$ $-\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}h_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}n_{j,-\beta},$ (26)
$T^{0}=$ $-\sum_{i,j}\sum_{\alpha\beta}\tilde{t}_{ij,\alpha\beta}\left(h_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}h_{j,-\beta}+n_{i,-\alpha}c^{\dagger}_{i\alpha}c_{j\beta}n_{j,-\beta}\right),$ (27)
@@
$S$ is calculated order by order in $\tilde{t}/U$ so that fluctuations to the doubly occupied sector are eliminated at each order in $\tilde{t}/U$ . We can expand the generator $S$, as $S=S^{(1)}+S^{(2)}+S^{(3)}+\cdots$, where $S^{n}\propto\left(\tilde{t}/U\right)^{n}$. To compute ring-exchange type terms we need to calculate $H_{\text{eff}}$ to $\mathcal{O}(\tilde{t}^{4}/U^{3})$. Therefore we require the expansion of the generator $S$ to $\mathcal{O}(\tilde{t}^{3}/U^{3})$ or up to $S^{(3)}$. We find that the first three terms of the expansion are:
@@
$iS^{(2)}=$ $\frac{1}{U^{2}}\left([T^{+},T^{0}]+[T^{-},T^{0}]\right),$ (32)
$iS^{(3)}=$ $\frac{1}{U^{3}}([[T^{+},T^{0}],T^{0}]-[[T^{-},T^{0}],T^{0}])$
@@ Appendix B Full model with Spin-Orbit interactions
$+$ $\sum_{(ijkl)\in R}J_{r_{2}}\left(\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{j}\right)\left(\mathbf{S}_{k}\overset{z}{\cdot}\mathbf{S}_{l}\right)+\left(\mathbf{S}_{j}\overset{z}{\cdot}\mathbf{S}_{k}\right)\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{l}\right)-2\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{k}\right)\left(\mathbf{S}_{j}\overset{z}{\cdot}\mathbf{S}_{l}\right)\right)$ (41)
$+$ $\sum_{(ijkl)\in R}J_{r_{2}}\left(\left(\mathbf{S}_{i}\cdot\mathbf{S}_{j}\right)\left(\mathbf{S}_{k}\cdot\mathbf{S}_{l}\right)+\left(\mathbf{S}_{j}\cdot\mathbf{S}_{k}\right)\left(\mathbf{S}_{i}\cdot\mathbf{S}_{l}\right)+2\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{k}\right)\left(\mathbf{S}_{j}\cdot\mathbf{S}_{l}\right)-2\left(\mathbf{S}_{i}\cdot\mathbf{S}_{k}\right)\left(\mathbf{S}_{j}\overset{z}{\cdot}\mathbf{S}_{l}\right)\right)$ (42)
@@
$+$ $\sum_{\langle ijkl\rangle}J_{r_{4}}\left(\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{j}\right)\left(\mathbf{S}_{k}\overset{z}{\cdot}\mathbf{S}_{l}\right)+\left(\mathbf{S}_{j}\overset{z}{\cdot}\mathbf{S}_{k}\right)\left(\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{l}\right)-\left(\mathbf{S}_{i}\cdot\mathbf{S}_{k}\right)\left(\mathbf{S}_{j}\cdot\mathbf{S}_{l}\right)\right),$ (44)
where the sum $(ijkl)\in R$ is over the three types of rhombuses in a triangular lattice, and $i$ is a sharp vertex of the rhombus $(i,j,k,l)$ counted anti-clockwise (see Fig. 1 (a)). We have defined the modified dot product as $\mathbf{S}_{i}\overset{z}{\cdot}\mathbf{S}_{j}=S^{z}_{i}S^{z}_{j}-S^{x}_{i}S^{x}_{j}-S^{y}_{i}S^{y}_{j}$. We obtain the strength of the exchange interactions from the strong coupling expansion in terms of the parameters, $t/U$ and $v_{z}/t$, of the original spin-orbit coupled Hubbard model.
@@ Appendix D DMRG Data
Using iDMRG we obtain a set of useful data that is used to characterize the nature of the quantum ground state. Firstly, iDMRG gives the ground state energy per lattice site. In addition to local magnetization $\langle S^{z}_{i}\rangle$, we obtain long-range real space spin correlations, $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle$, along the infinite direction $x$ of the cylinder. From the real space spin correlations, we can calculate the static spin structure factor (10). We also obtain real space long-ranged dimer correlations of the spins, $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle$, where $D^{n}_{i}=\mathbf{S}_{i}\cdot\mathbf{S}_{i+\mathbf{a}_{n}}$ is the dimer operator along the direction $\mathbf{a}_{n}$ (see Fig. 1 (a)). Using this the static dimer structure factor (12) can be calculated. Further, we also obtain the scalar chirality $\chi$ as defined in (9), and the handedness $\omega$ as defined in (11). Most importantly, from iDMRG we can obtain the momentum-resolved entanglement spectrum. Figs. 9-13 show the various correlation functions and entanglement spectra for all the different phases appearing in Figs. 2 & 3.
@@
Figure 9: Observables in the CSL phase ($J_{1}=1$, $J_{2}=0.01$, $J_{3}=0.01$, $J_{r}=0.2$, $D_{z}/J_{1}=0.04$, $J_{r_{1}}/J_{r}=0.4$). The real space spin correlations $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle_{c}=\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle$ are shown in (a), the static spin structure factor $S(\mathbf{k})$ in (b). Real space dimer correlations $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle_{c}=\left\langle D^{n}_{i}D^{n}_{j}\right\rangle-\left\langle D^{n}_{i}\right\rangle\left\langle D^{n}_{j}\right\rangle$ along the the different directions $\mathbf{a}_{n}$ are shown in (d)-(f), and the static dimer structure factors $D_{n}(\mathbf{k})$ in (g)-(i). Here we have chosen $\mathbf{R}_{0}=3\mathbf{a}_{1}$. The entanglement spectrum in (c) shows the characteristic structure of the Kalmeyer-Laughlin type CSL [68, 70].
@@
Figure 13: Observables in the magnetic ordered (MO) phase ($J_{1}=1$, $J_{2}=0.012$, $J_{3}=0.012$, $J_{r}=0.24$, $D_{z}=0.065$, $J_{r_{1}}=0.072$). The real space spin correlations $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle_{c}=\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle$ are shown in (a), the static spin structure factor $S(\mathbf{k})$ in (b). Real space dimer correlations $\left\langle D^{n}_{i}D^{n}_{j}\right\rangle_{c}=\left\langle D^{n}_{i}D^{n}_{j}\right\rangle-\left\langle D^{n}_{i}\right\rangle\left\langle D^{n}_{j}\right\rangle$ along the the different directions $\mathbf{a}_{n}$ are shown in (d)-(f), and the static dimer structure factors $D_{n}(\mathbf{k})$ in (g)-(i). Here we have chosen $\mathbf{R}_{0}=3\mathbf{a}_{1}$. Even though the spin correlations in (a) are long-ranged, they are shorter-ranged than the $120^{\circ}_{\pm}$ phases in Figs. 11, 12.
@@ Appendix E Different types of $120^{\circ}$ order
[…]
@@
[…]